# On the Mode-Seeking Properties of Langevin Dynamics

### v2

@@ 3.2 Multi-Modal Distributions
Our work focuses on multi-modal distributions. We use $P=\sum_{i\in[k]}w_{i}P^{(i)}$ to represent a mixture of $k$ modes, where each mode $P^{(i)}$ is a probability density with frequency $w_{i}$ such that $w_{i}>0$ for all $i\in[k]$ and $\sum_{i\in[k]}w_{i}=1$. In our theoretical analysis, we consider Gaussian mixtures and sub-Gaussian mixtures, i.e., every component $P^{(i)}$ is a Gaussian or sub-Gaussian distribution. A probability distribution $p({\mathbf{z}})$ of dimension $d$ is defined as a sub-Gaussian distribution with parameter ${\nu}^{2}$ if, given the mean vector ${\bm{\mu}}:=\mathbb{E}_{{\mathbf{z}}\sim p}[{\mathbf{z}}]$, the moment generating function (MGF) of $p$ satisfies the following inequality for every vector ${\bm{\alpha}}\in{\mathbb{R}}^{d}$:
We remark that sub-Gaussian distributions include a wide variety of distributions such as Gaussian distributions and any distribution within a bounded $\ell_{2}$-norm distance from the mean ${\bm{\mu}}$.
4 Theoretical Analysis of the Mode-Seeking Properties of Langevin Dynamics
In this section, we theoretically investigate the mode-seeking properties of vanilla Langevin dynamics. We begin with analyzing Langevin dynamics in Gaussian mixtures, and further generalize the results to sub-Gaussian mixtures. We again highlight that in our theoretical analysis, we assume the sampler has access to the score function $\nabla_{\mathbf{x}}\log P({\mathbf{x}})$ of the underlying distribution $P$.
4.1 Langevin Dynamics in Gaussian Mixtures
Assumption 1.
Consider a data distribution $P:=\sum_{i=0}^{k}w_{i}P^{(i)}$ as a mixture of Gaussian distributions, where $1\leq k=o(d)$ and $w_{i}>0$ is a positive constant such that $\sum_{i=0}^{k}w_{i}=1$. Suppose that $P^{(i)}={\mathcal{N}}({\bm{\mu}}_{i},{\nu}_{i}^{2}{\bm{I}}_{d})$ is a Gaussian distribution over ${\mathbb{R}}^{d}$ for all $i\in\{0\}\cup[k]$ such that for all $i\in[k]$, ${\nu}_{i}<{\nu}_{0}$ and $\left\|{{\bm{\mu}}_{i}-{\bm{\mu}}_{0}}\right\|^{2}\leq\frac{{\nu}_{0}^{2}-{\nu}_{i}^{2}}{2}\left(\log\left(\frac{{\nu}_{i}^{2}}{{\nu}_{0}^{2}}\right)-\frac{{\nu}_{i}^{2}}{2{\nu}_{0}^{2}}+\frac{{\nu}_{0}^{2}}{2{\nu}_{i}^{2}}\right)d$. Denote ${\nu}_{\max}:=\max_{i\in[k]}{\nu}_{i}$.
To intuitively understand Assumption 1, we first note that the probability density $p({\mathbf{z}})$ of a Gaussian distribution ${\mathcal{N}}({\bm{\mu}},{\nu}^{2}{\bm{I}}_{d})$ decays exponentially in terms of $\frac{\left\|{{\mathbf{z}}-{\bm{\mu}}}\right\|^{2}}{{\nu}^{2}}$. When a state ${\mathbf{z}}$ is sufficiently far from one mode $P^{(i)}$, the probability density of $P^{(i)}$ is dominated by the high-variance component $P^{(0)}$, which implies that the gradient information from $P^{(i)}$ will be masked by $P^{(0)}$. Hence, the dynamics can only visit the universal mode unless the stochastic noise miraculously leads it to the region of another mode. In addition, it can be verified that $\log\left(\frac{{\nu}_{i}^{2}}{{\nu}_{0}^{2}}\right)-\frac{{\nu}_{i}^{2}}{2{\nu}_{0}^{2}}+\frac{{\nu}_{0}^{2}}{2{\nu}_{i}^{2}}$ is a positive constant for ${\nu}_{i}<{\nu}_{0}$, thus the last requirement of Assumption 1 essentially represents $\left\|{{\bm{\mu}}_{i}-{\bm{\mu}}_{0}}\right\|^{2}\leq{\mathcal{O}}(d)$. We formalize the intuition in Theorem 1 and defer the proof to Appendix A.1.
@@ 4.1 Langevin Dynamics in Gaussian Mixtures
Consider a data distribution $P$ satisfying Assumption 1. We follow Langevin dynamics for $T=\exp({\mathcal{O}}(d))$ steps. Suppose the sample is initialized in $P^{(0)}$, then with probability at least $1-T\cdot\exp(-\Omega(d))$, we have $\left\|{{\mathbf{x}}_{t}-{\bm{\mu}}_{i}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$ for all $t\in\{0\}\cup[T]$ and $i\in[k]$.
The constants in the notation $\Omega(d)$ are specified in Equations 6 and 7 in the Appendix. We note that $\left\|{{\mathbf{x}}_{t}-{\bm{\mu}}_{i}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$ is a strong notion of mode-seekingness, since the density of mode $P^{(i)}={\mathcal{N}}({\bm{\mu}}_{i},{\nu}_{i}^{2}{\bm{I}}_{d})$ concentrates around the $\ell_{2}$-norm ball $\left\{{{\mathbf{z}}:\left\|{{\mathbf{z}}-{\bm{\mu}}_{i}}\right\|^{2}\leq{\nu}_{i}^{2}d}\right\}$. This notion can be translated into a lower bound in terms of other distance measures, e.g., total variation distance in the following corollary, whose proof is deferred to Appendix A.2.
@@
Under the same assumptions as in Theorem 1, for all time steps $t\in\{0\}\cup[T]$, the distribution $\hat{P}_{t}$ of the generated sample ${\mathbf{x}}_{t}$ by Langevin dynamics at time $t$ satisfies
$\text{TV}(\hat{P}_{t},P)\geq(1-w_{0})\left(1-\frac{T}{\exp(\Omega(d))}\right).$
4.2 Langevin Dynamics in Sub-Gaussian Mixtures
We further generalize our results to sub-Gaussian mixtures. We impose the following assumptions on the mixture. It is worth noting that Assumptions 2.i.-iii. automatically hold for Gaussian mixtures such that $P^{(i)}={\mathcal{N}}({\bm{\mu}}_{i},{\nu}_{i}^{2}I)$, and Assumptions 2.iv. and v. are specific assumptions on the mean and variance of $P^{(i)}$, similar to Assumption 1.
Assumption 2.
  1. i.
  2. ii.
    $P^{(i)}$ is differentiable and $\nabla P^{(i)}({\bm{\mu}}_{i})=\bm{0}_{d}$,
  3. iii.
  4. iv.
  5. v.
The feasibility of Assumption 2.v. is validated by Lemma 9 in Appendix A.3. With Assumption 2, we show the mode-seeking tendency of Langevin dynamics under sub-Gaussian distributions in Theorem 2 and defer the proof to Appendix A.3.
@@ 4.2 Langevin Dynamics in Sub-Gaussian Mixtures
Consider a data distribution $P$ satisfying Assumption 2. We follow Langevin dynamics for $T=\exp({\mathcal{O}}(d))$ steps. Suppose the sample is initialized in $P^{(0)}$, then with probability at least $1-T\cdot\exp(-{\mathcal{O}}(d))$, we have $\left\|{{\mathbf{x}}_{t}-{\bm{\mu}}_{i}}\right\|^{2}>\left(\frac{{\nu}_{0}^{2}}{2}+\frac{{\nu}_{\max}^{2}}{2(1-c_{{\nu}})}\right)d$ for all $t\in\{0\}\cup[T]$ and $i\in[k]$.
We remark that an implication of Theorem 2 is the potential difficulty of transition between low-variance modes for Langevin dynamics. For instance, suppose $P^{(1)},\cdots,P^{(k)}$ have bounded support sets with small radius and sufficiently distant means. If the sample is initialized in a low-variance mode $P^{(m)}$ (for $m\in[k]$), either it stays in $P^{(m)}$ and cannot capture other modes, or it escapes $P^{(m)}$ (due to the random noise $\bm{\epsilon}$) and is expected to need to explore the whole space until finding the support sets of the other bounded modes.
Furthermore, in Appendix B we extend our theoretical analysis to annealed Langevin dynamics with bounded noise levels, indicating the effect of annealing noise levels on the mode-seeking tendencies of Langevin dynamics. Aligning with the empirical analysis in [26], we show that bounded noise levels will have a limited impact on Langevin dynamics since they exhibit similar mode-seeking tendencies. On the other hand, as suggested by [26], annealed Langevin dynamics with a significantly larger initial noise level could capture more modes, which is also confirmed by our numerical results in Section 6.
@@ 5 Chained Langevin Dynamics
To reduce the mode-seeking tendencies of vanilla Langevin dynamics, we propose Chained Langevin Dynamics (Chained-LD) in Algorithm 1. While vanilla Langevin dynamics apply gradient updates to all coordinates of the sample in every step, we decompose the sample into patches of constant size and generate each patch sequentially to alleviate the exponential dependency on the dimensionality. More precisely, we divide a sample ${\mathbf{x}}$ into $d/Q$ patches ${\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(d/Q)}$ of some constant size $Q$, and apply Langevin dynamics to sample each patch ${\mathbf{x}}^{(q)}$ (for $q\in[d/Q]$) from the conditional distribution $P({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(q-1)})$. Intuitively, vanilla Langevin dynamics needs to explore the entire space (of volume exponentially large in $d$) to find the missing modes, while Chained-LD could significantly lower the volume by dimensionality reduction.
@@
1:Patch size $Q$, dimension $d$, conditional score function $\nabla\log P_{{\sigma}_{t}}$, number of iterations $T$, noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[TQ/d]}$, step size $\left\{{\delta_{t}}\right\}_{t\in[TQ/d]}$.
@@
In practice, we can also apply annealed Langevin dynamics [4] to facilitate the sampling of each patch, by perturbing it with a series of noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[TQ/d]}$. Specifically, we refer chained vanilla Langevin dynamics (Chained-VLD) to Algorithm 1 without noise injection (i.e., ${\sigma}_{t}=0$ for all $t\in[TQ/d]$), and chained annealed Langevin dynamics (Chained-ALD) otherwise. Ideally, if a sampler perfectly generates every patch, combining all patches gives a sample from the original distribution due to the chain rule $P({\mathbf{x}})=\prod_{q\in[d/Q]}P({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(q-1)})$. In Proposition 1 we give a linear reduction from producing samples of dimension $d$ using Chained-LD to learning the distribution of a $Q$-dimensional variable for constant $Q$. The proof of Proposition 1 is deferred to Appendix C.
Proposition 1.
Consider a sampler algorithm taking the first $q-1$ patches ${\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(q-1)}$ as input and outputing a sample of the next patch ${\mathbf{x}}^{(q)}$ with probability $\hat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(q-1)}\right)$ for all $q\in[d/Q]$. Suppose that for every $q\in[d/Q]$ and any given previous patches ${\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(q-1)}$, the sampler algorithm can achieve
$\text{TV}\left(\hat{P}\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(q-1)}\right),P\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(q-1)}\right)\right)\leq\varepsilon\cdot\frac{Q}{d}$
in $\tau(\varepsilon/d)$ iterations for some $\varepsilon>0$. Then, equipped with the sampler algorithm, the Chained-LD algorithm in $\frac{d}{Q}\cdot\tau(\varepsilon/d)$ iterations can achieve
$\text{TV}\left(\hat{P}({\mathbf{x}}),P({\mathbf{x}})\right)\leq\varepsilon.$
With additional assumptions on the target distribution $P$, we can obtain upper bounds on the iteration complexity of Chained-LD. If the conditional distribution $p:=P\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(q-1)}\right)$ for all $q\in[d/Q]$ satisfies the assumptions specified in Appendix A of [37], i.e., $\log p$ is $L$-Lipschitz smooth for all ${\mathbf{x}}^{(q)}$ and $m$-strongly concave for ${\mathbf{x}}^{(q)}$ outside an $\ell_{2}$-norm ball of radius $R$ (i.e., $\left\|{{\mathbf{x}}^{(q)}}\right\|>R$), by Theorem 1 of [37] we obtain $\tau(\varepsilon/d)={\mathcal{O}}\left(\exp(32LR^{2})\frac{L^{2}}{m^{2}}\cdot\frac{R}{(\varepsilon/d)^{2}}\log\frac{R}{(\varepsilon/d)^{2}}\right)$. In common settings, the radius $R$ in $Q$–dimensional space scales as $R=\Theta(\sqrt{Q})$. Therefore, for constants $L$ and $m$, the iteration complexity of Chained-LD is a polynomial ${\mathcal{O}}\left(\exp({\mathcal{O}}(LQ))\frac{d^{3}}{\varepsilon^{2}\sqrt{Q}}\log\frac{d^{2}\sqrt{Q}}{\varepsilon^{2}}\right)$ in $d$, which takes its minimum value ${\mathcal{O}}\left(\frac{d^{3}}{\varepsilon^{2}}\log\frac{d^{2}}{\varepsilon^{2}}\right)$ when the patch size $Q$ is a constant not growing with $d$. For comparison, the iteration complexity of vanilla Langevin dynamics could scale exponentially in $d$ when $R=\Theta(\sqrt{d})$ in common settings.
Remark 2.
We highlight that Chained-LD is a sampling algorithm. In Algorithm 1, we assume the sampler has direct access to the conditional score function. Also, the conditional densities used in Chained-LD do not require any extra information compared to the target distribution assumed in vanilla LD, since vanilla LD has access to the joint distribution and Chained-LD has access to conditional distributions that, based on chain rule, have the same information as the joint distribution.
Remark 3.
We note that similar combinations of autoregressive models and denoising diffusion models have been studied in the generative modeling literature, in the context of text generation [38, 39] and time series forecasting [40]. At a high level, the generative modeling literature focuses on the training and implementation of autoregressive diffusion models in time-dependent scenarios, while this work focuses on the theoretical guarantees of Chained Langevin dynamics and its comparison with vanilla Langevin dynamics, motivated by their mode-seeking properties under sub-Gaussian mixture distribution.

### v3

@@ 3.2 Multi-Modal Distributions
In this work, we focus on the analysis of Langevin dynamics in multi-modal distributions. We highlight that our work studies Langevin dynamics under multi-modal distributions in a slightly different setting from the standard theory literature on sampling. The existing theoretical literature commonly considers a mixture of well-separated modes with bounded variance. On the other hand, in our analysis, we consider a low-density high-variance mode (referred to as Mode 0 or $P^{(0)}$) surrounding the other modes and filling the low density region between the modes. Specifically, as illustrated in Figure 2.(a), we formulate a symmetric 3-Gaussian model as a hard example for Langevin dynamics, defined as following
Definition 1.
For any given frequency $w\in(0,1)$ and variance $\nu^{2}>1$ of the in-between mode $P^{(0)}$, and any mean vector ${\bm{\mu}}$ of the low-variance mode $P^{(1)}$, a symmetric 3-Gaussian model is defined as
$P_{w,\nu,{\bm{\mu}}}=w{\mathcal{N}}(\bm{0},\nu^{2}{\bm{I}})+\frac{1-w}{2}{\mathcal{N}}({\bm{\mu}},{\bm{I}})+\frac{1-w}{2}{\mathcal{N}}(-{\bm{\mu}},{\bm{I}}).$
More generally, we use $P=w_{0}P^{(0)}+\sum_{i\in[k]}w_{i}P^{(i)}$ to represent a mixture of $k+1$ modes, where $P^{(0)}$ is the in-between mode with high variance as illustrated in Figure 2.(b). Here each mode $P^{(i)}$ is a probability density with frequency $w_{i}$ such that $w_{i}>0$ for all $i\in[k]$ and $w_{0}+\sum_{i\in[k]}w_{i}=1$.
4 Theoretical Analysis of the Hardness of Langevin Dynamics
In this section, we theoretically investigate the iteration complexity of Langevin dynamics. We first introduce a notation $\left\|{{\mathbf{x}}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}$ to measure the distance between a sample ${\mathbf{x}}\in{\mathbb{R}}^{d}$ to the linear span of mean vectors $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$ of the mixture components in a multi-modal distribution.
Definition 2.
For a sample ${\mathbf{x}}\in{\mathbb{R}}^{d}$ and a set of vectors ${\bm{\mu}}_{1},\cdots,{\bm{\mu}}_{k}\in{\mathbb{R}}^{d}$, we define $\left\|{{\mathbf{x}}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}$ as the distance from ${\mathbf{x}}$ to the span of $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$, i.e., the minimum distance from ${\mathbf{x}}$ to any linear combination of $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$:
$\left\|{{\mathbf{x}}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}:=\min_{\lambda_{1},\cdots,\lambda_{k}}\big{\|}{{\mathbf{x}}-\sum_{i=1}^{k}\lambda_{i}{\bm{\mu}}_{i}}\big{\|}.$ (4)
We aim to show that in a mixture distribution $P$ with a high-variance mode $P^{(0)}$, the sampled vector ${\mathbf{x}}$ is likely to be far from the low-variance modes $P^{(1)},\cdots,P^{(k)}$ in terms of the $\left\|{{\mathbf{x}}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}$ metric.
4.1 Langevin Dynamics in Symmetric 3-Gaussian Model
We begin our theoretical analysis with a simple case: a symmetric 3-Gaussian model consisting of two symmetric Gaussian modes $P^{(1)}={\mathcal{N}}({\bm{\mu}},{\bm{I}}_{d})$ and $P^{(2)}={\mathcal{N}}(-{\bm{\mu}},{\bm{I}}_{d})$, and an in-between mode $P^{(0)}={\mathcal{N}}(\bm{0}_{d},\nu^{2}{\bm{I}}_{d})$ with high variance $\nu^{2}\geq 3$, as illustrated in Figure 2.(a). In the following Theorem 1, we show that with high probability, the sampled vector ${\mathbf{x}}_{T}$ fails to find the symmetric modes $P^{(1)},P^{(2)}$ within a sub-exponential number of iterations. The proof of Theorem 1 is deferred to Appendix A.1.
@@ 4.1 Langevin Dynamics in Symmetric 3-Gaussian Model
Consider a distribution $P_{w,\nu,{\bm{\mu}}}=w{\mathcal{N}}(\bm{0}_{d},\nu^{2}{\bm{I}}_{d})+\frac{1-w}{2}{\mathcal{N}}({\bm{\mu}},{\bm{I}}_{d})+\frac{1-w}{2}{\mathcal{N}}(-{\bm{\mu}},{\bm{I}}_{d})$ by Definition 1 in dimension $d\geq 250$, such that $w\geq 0.01$, $\nu^{2}\geq 3$, and $\left\|{{\bm{\mu}}}\right\|^{2}\leq 0.2d$. We initialize the sample ${\mathbf{x}}_{0}$ such that $\left\|{{\mathbf{x}}_{0}}\right\|_{{\bm{\mu}}}^{2}\geq\frac{3\nu^{2}+1}{4}d$ and apply Langevin dynamics for $T$ iterations, then we have
${\mathbb{P}}\left(\left\|{{\mathbf{x}}_{T}}\right\|_{{\bm{\mu}}}^{2}\geq\frac{\nu^{2}+1}{2}d\right)\geq 1-T\cdot\exp\left(-\frac{d}{300}\right).$
For example, for a symmetric 3-Gaussian model $P_{0.01,\sqrt{3},0.2\cdot\bm{1}_{d}}$, Theorem 1 indicates that the sampled vector ${\mathbf{x}}_{T}$ within $T\leq\exp(d/300)$ iterations cannot be $\sqrt{2d}$ close to the center of any low-variance modes with high probability. To interpret Theorem 1, we first note that in a high-dimensional space ${\mathbb{R}}^{d}$, the probability mass of a Gaussian distribution ${\mathcal{N}}({\bm{\mu}},{\bm{I}}_{d})$ concentrates inside a ball of radius $\sqrt{d}$ centered at ${\bm{\mu}}$, i.e., $\left\|{{\mathbf{x}}-{\bm{\mu}}}\right\|^{2}\leq d$. On the other hand, the high probability bound $\left\|{{\mathbf{x}}_{T}}\right\|_{{\bm{\mu}}}^{2}\geq\frac{\nu^{2}+1}{2}d$ in Theorem 1 implies that ${\mathbf{x}}_{T}$ is far from the center of both symmetric Gaussian modes, i.e., $\left\|{{\mathbf{x}}_{T}-{\bm{\mu}}}\right\|^{2}\geq\frac{\nu^{2}+1}{2}d\geq 2d$. This observation allows us to translate the bound on $\left\|{{\mathbf{x}}_{T}}\right\|_{{\bm{\mu}}}$ into a lower bound in other standard metrics such as total variation distance, as shown in the following Corollary 1.
@@
Under the same assumptions as in Theorem 1, the distribution $\widehat{P}_{T}$ of the sampled vector ${\mathbf{x}}_{T}$ by Langevin dynamics satisfies
$\text{TV}(\widehat{P}_{T},P)\geq 0.99-w-\frac{T}{\exp(-d/300)}.$
4.2 Langevin Dynamics in Gaussian Mixture Models
We further extend Theorem 1 to a general Gaussian mixture setting. As illustrated in Figure 2.(b), we consider a Gaussian mixture with an in-between mode $P^{(0)}$ with high variance. To intuitively understand our theoretical results, we first note that the probability density $p({\mathbf{z}})$ of a Gaussian distribution ${\mathcal{N}}({\bm{\mu}},\nu^{2}{\bm{I}}_{d})$ decays exponentially in terms of $\frac{\left\|{{\mathbf{z}}-{\bm{\mu}}}\right\|^{2}}{\nu^{2}}$. When a sample ${\mathbf{z}}$ is sufficiently far from one mode $P^{(i)}$, since $P^{(0)}$ has a higher variance, the probability density of $P^{(i)}$ is dominated by mode $P^{(0)}$ and the gradient information from $P^{(i)}$ will be masked by $P^{(0)}$. Hence, the dynamics can only visit $P^{(0)}$ unless the stochastic noise miraculously leads it to the region of another low-variance mode.We formalize this intuition in Theorem 2 and defer the proof to Appendix A.2.
@@ 4.2 Langevin Dynamics in Gaussian Mixture Models
Consider a data distribution $P=w_{0}{\mathcal{N}}(\bm{0}_{d},{\nu}_{0}^{2}{\bm{I}}_{d})+\sum_{i\in[k]}w_{i}{\mathcal{N}}({\bm{\mu}}_{i},{\nu}_{i}^{2}{\bm{I}}_{d})$ in dimension $d\geq 250$. For all low-variance modes $P^{(1)},\cdots,P^{(k)}$, we assume $\left\|{{\bm{\mu}}_{i}}\right\|\leq 0.2d$ and denote ${\nu}_{\max}:=\max_{i\in[k]}{\nu}_{i}$. For in-between mode $P^{(0)}$, assume $w_{0}\geq 0.01$ and ${\nu}_{0}^{2}\geq 3{\nu}_{\max}^{2}$. We initialize the sample ${\mathbf{x}}_{0}$ such that $\left\|{{\mathbf{x}}_{0}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$ and apply Langevin dynamics for $T$ iterations, then we have
${\mathbb{P}}\left(\left\|{{\mathbf{x}}_{T}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}\geq\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d\right)\geq 1-T\cdot\exp\left(-\frac{d}{300}\right).$
4.3 Iteration Complexity of Annealed Langevin Dynamics
Next, we generalize our theoretical results to annealed Langevin dynamics with bounded noise levels in Theorem 3, under similar assumptions on the target distribution. The proof is deferred to Appendix A.3. Aligning with the analysis in (song2020improved, ), we show that bounded noise levels will have a limited impact on Langevin dynamics since they exhibit similar exponential complexity in high-dimensional distributions. On the other hand, as suggested by (song2020improved, ), annealed Langevin dynamics with a significantly larger initial noise level ${\sigma}_{0}$ could capture more modes (e.g., ${\sigma}_{0}={\mathcal{O}}(\sqrt{d})$), which is also confirmed by our numerical results in Section 6.
Theorem 3.
Consider a data distribution $P=w_{0}{\mathcal{N}}(\bm{0}_{d},{\nu}_{0}^{2}{\bm{I}}_{d})+\sum_{i\in[k]}w_{i}{\mathcal{N}}({\bm{\mu}}_{i},{\nu}_{i}^{2}{\bm{I}}_{d})$ in dimension $d\geq 250$. For all low-variance modes $P^{(1)},\cdots,P^{(k)}$, we assume $\left\|{{\bm{\mu}}_{i}}\right\|\leq 0.05d$ and denote ${\nu}_{\max}:=\max_{i\in[k]}{\nu}_{i}$. For in-between mode $P^{(0)}$, assume $w_{0}\geq 0.01$ and ${\nu}_{0}^{2}\geq 3{\nu}_{\max}^{2}$. We initialize the sample ${\mathbf{x}}_{0}$ such that $\left\|{{\mathbf{x}}_{0}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d+{\sigma}_{0}^{2}d$ and apply annealed Langevin dynamics for $T$ steps with noise levels ${\nu}_{\max}\geq{\sigma}_{0}\geq\cdots\geq{\sigma}_{T}\geq 0$, then we have
${\mathbb{P}}\left(\left\|{{\mathbf{x}}_{0}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}^{2}\geq\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d\right)\geq 1-T\cdot\exp\left(-\frac{d}{1500}\right).$
Finally, in Appendix B, we extend our theoretical results to sub-Gaussian mixtures $P=w_{0}P^{(0)}+\sum_{i\in[k]}w_{i}P^{(i)}$, where $P^{(i)}$ is a sub-Gaussian distribution of mean ${\bm{\mu}}_{i}$ with parameter ${\nu}_{i}^{2}$ satisfying that the score function of $P^{(i)}$ is Lipschitz. We show that if the sample ${\mathbf{x}}_{0}$ is initialized far from the mean vectors, Langevin dynamics and annealed Langevin dynamics still exhibit similar exponential complexity to converge to low-variance sub-Gaussian modes in the target distribution.
@@ 5 Chained Langevin Dynamics
To reduce the exponential complexity of Langevin dynamics, we propose Chained Langevin Dynamics (Chained-LD) in Algorithm 1. While Langevin dynamics apply gradient updates to all coordinates of the variable at every step, we decompose the variable into patches of constant size and sample each patch sequentially to alleviate the exponential dependency on the dimensionality. More precisely, we divide a vector ${\mathbf{x}}$ into $d/Q$ patches ${\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(d/Q)}$ of some constant size $Q$, and apply Langevin dynamics to sample each patch ${\mathbf{x}}^{(q)}$ (for $q\in[d/Q]$) from the conditional distribution $P({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(q-1)})$. Intuitively, vanilla Langevin dynamics needs to explore the entire space (of volume exponentially large in $d$) to find the missing modes, while Chained-LD could significantly lower the volume by dimensionality reduction.
@@
1:Patch size $Q$, dimension $d$, number of iterations $T$, noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[TQ/d]}$, conditional score function $\nabla\log P_{{\sigma}_{t}}$, step size $\left\{{\delta_{t}}\right\}_{t\in[TQ/d]}$.
@@
We can also apply annealed Langevin dynamics (song2019generative, ) to facilitate the sampling of each patch, by perturbing it with a series of noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[TQ/d]}$. Specifically, we refer chained vanilla Langevin dynamics (Chained-VLD) to Algorithm 1 without noise injection (i.e., ${\sigma}_{t}=0$ for all $t\in[TQ/d]$), and chained annealed Langevin dynamics (Chained-ALD) otherwise. Ideally, if a sampler perfectly generates every patch, combining all patches gives a vector from the original distribution due to the chain rule
$P({\mathbf{x}})=\prod_{q\in[d/Q]}P({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(q-1)}).$
In Theorem 4, we prove that Chained-LD can provably converge to the target distribution within $\varepsilon$ total variation distance, in a polynomial number of iterations. Similar to (cheng2018sharp, ; ma2019sampling, ), we assume that the log conditional PDF of every patch $\log P({\mathbf{x}}^{(q)}|{\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(q-1)})$ is $L_{Q}$-smooth and $m_{Q}$-strongly concave for ${\mathbf{x}}^{(q)}>R_{Q}$. The details of Assumption 4 and the proof of Theorem 4 is deferred to Appendix C.
Theorem 4.
Consider a data distribution $P$ satisfying Assumption 4. We initialize ${\mathbf{x}}_{0}\sim{\mathcal{N}}(\bm{0}_{d},\frac{1}{L_{Q}}{\bm{I}}_{d})$ and apply chained Langevin dynamics in Algorithm 1 with constant patch size $Q$, noise level ${\sigma}_{t}=0$, and step size $\delta_{t}=\frac{m_{Q}\varepsilon^{2}Q}{64L_{Q}^{2}d^{2}}\exp(-16L_{Q}R_{Q}^{2})$. Then, for
$T=\frac{128L_{Q}^{2}d^{3}}{m_{Q}^{2}Q^{2}\varepsilon^{2}}\exp(32L_{Q}R_{Q}^{2})\log\left(\frac{d^{3}}{\varepsilon^{2}Q^{2}}\right),$
the output distribution $\widehat{P}({\mathbf{x}})$ after $T$ iterations satisfies $\text{TV}(\widehat{P}({\mathbf{x}}),P({\mathbf{x}}))\leq\varepsilon$ for any constant $\varepsilon>0$.
We highlight that due to dimension reduction, in general, the parameters $L_{Q},m_{Q},R_{Q}$ are constants that do not grow with dimension $d$. To give a concrete example, we consider a symmetric 3-Gaussian model
$P_{w,\nu,\bm{1}_{d}}=w{\mathcal{N}}(\bm{0}_{d},\nu^{2}{\bm{I}}_{d})+\frac{1-w}{2}{\mathcal{N}}(\bm{1}_{d},{\bm{I}}_{d})+\frac{1-w}{2}{\mathcal{N}}(-\bm{1}_{d},{\bm{I}}_{d}).$
Then, for every patch $q\in[d/Q]$, the conditional distribution is given by
$P\left({\mathbf{x}}^{(q)}|{\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(q-1)}\right)=w{\mathcal{N}}(\bm{0}_{Q},\nu^{2}{\bm{I}}_{Q})+\frac{1-w}{2}{\mathcal{N}}(\bm{1}_{Q},{\bm{I}}_{d})+\frac{1-w}{2}{\mathcal{N}}(-\bm{1}_{Q},{\bm{I}}_{Q}),$
which is independent from the dimension $d$ of the whole vector ${\mathbf{x}}$. Therefore, the parameters $L_{Q},m_{Q},R_{Q}$ depend only on the patch size $Q$, which is set as a constant. In contrast, without dimension reduction, $-\log P_{w,\nu,\bm{1}_{d}}({\mathbf{x}})$ is non-convex for ${\mathbf{x}}=\bm{1}_{d}$. Therefore, under the assumption that the distribution of the whole vector $-\log P_{w,\nu,\bm{1}_{d}}({\mathbf{x}})$ is strongly-convex for $\left\|{{\mathbf{x}}}\right\|>R$ where $R>\sqrt{d}$, the upper bound on the iteration complexity of Langevin dynamics obtained by (cheng2018sharp, ) and (ma2019sampling, ) scales as ${\mathcal{O}}(\exp(cLR^{2}\text{poly}(d,1/\varepsilon)))\geq{\mathcal{O}}(\exp(cLd))$, which is exponential in dimension $d$.